OpenStudy (anonymous):

If you were to use the substitution method to solve the following system, choose the new system of equations that would result if x was isolated in the second equation. 2x + 5y – 2z = –1 x + 3y + z = –2 3x – 2y + 3z = –17 –y – 4z = 3 –11y = –11 11y = 3 7y + 6z = –11 4x + 11y = –5 –11y = –11 –y = –5 6x + 7y = –11

OpenStudy (radar):

The second equation: x + 3y + z= -2 x=-3y -x - 2 x is now isolated

OpenStudy (radar):

We now substitute for x in the first and third equation and see what you get.

OpenStudy (anonymous):

its the 2nd one?

OpenStudy (radar):

Let me check. 2(-3y-z-2) + 5y-2z=-1 note typo in y first post change -x to -z in isolations of x. -6y-2z-4 + 5y-2z=-1 -y-4z=3 yes the second one as the remaining does not have that so no need to verify the third equation.

OpenStudy (anonymous):

thanks so much

OpenStudy (radar):

No problem for my own edification I will also do the third one to verify. 3(-3y-z-2)-2y+3z=-17 -9y-3z-6-2y+3z=-17 -11y=-11 Yep.

OpenStudy (anonymous):

can you help me with the other one i just posted